OpenStudy (loser66):

Use the differentiation operator on \(P_n\) to show that the range and the null space of a linear transformation need Not be disjoint Please help.

OpenStudy (helder_edwin):

is \(P_n\) the set of polynomials of degree <= n ?

OpenStudy (loser66):

yes, please

OpenStudy (helder_edwin):

first of all \[\large P_n=\langle1,x,x^2,\dots,x^n\rangle. \] if \(D:P_n\to P_n\) is given by \(D(p(x))=d p(x)/dx\), then \[\large \ker D=\langle 1\rangle \] and \[\large \text{im}\ D=\langle1,2x,3x^2,\dots,nx^{n-1}\rangle= \langle1,x,x^2,\dots,x^{n-1}\rangle \] therefore \[\large \ker D\cap\text{im }D=\langle1\rangle\neq\emptyset \]

OpenStudy (loser66):

what does im (D) mean? I don't understand why im (D) = <1,2x,3x^2,....,nx^n-1> = ..... how and why? if it stop there, so I understand that is the set of target after take derivative, but you let it = < basis of P_n> it confused me.

OpenStudy (helder_edwin):

i don't know how u call it in english. "im" stands for image (imagen in spanish).

OpenStudy (loser66):

oh, you mean range (D)

OpenStudy (helder_edwin):

yes

OpenStudy (helder_edwin):

the second line P_n=<...> is the cannonical base of Pn.

OpenStudy (loser66):

how they are equal?

OpenStudy (helder_edwin):

i see. i guess u use the notation "span" so \[\large \langle1,x,\dots,x^n\rangle=\text{span}\{1,x,\dots,x^n\} \]

OpenStudy (loser66):

<1,x,......x^n> is basis of P_n (not span) but how it is = range (D)

OpenStudy (loser66):

oh, I got what you mean now. Thank a ton

OpenStudy (helder_edwin):

sure. u r welcome